# Which Optimizer Works Best for Physics-Informed Neural Networks and Kolmogorov-Arnold Networks?

### v4

@@ 2.1.2 Performance of optimzers based on composition operator in the JAX ecosystem
Figure 8: Convergence of PINNs for the viscous Burgers equation under various combinations of line search and BFGS algorithms for single-precision arithmetic (32-bit floats) with GPU fully saturated with model parameters to avoid the effect of latency on compute time: The left panel illustrates a scenario where the iteration count of the first-order Adam algorithm dominates, combined with BFGS using the trust-region approach. This pairing exhibits degeneracy due to limited numerical precision during BFGS iterations. The right panel presents the convergence behavior of BFGS paired with trust-region and backtracking line search methods. In this setup, Adam is applied for 11 iterations as a warmup. The BFGS-based optimizer requires significantly more iterations than first-order optimization methods like Adam and Rectified Adam (RAdam) [79].
Adam [16001] + BFGS with trust-region [13] $2.91\times 10^{-3}$ 127 3501
Adam [11] + BFGS with backtracking [2113] $7.25\times 10^{-4}$ 61 3501
RAdam [11] + BFGS with backtracking [1008] $7.67\times 10^{-4}$ 62 3501
BFGS with backtracking [1677] $2.48\times 10^{-3}$ 45 3501
BFGS with trust-region [944] $1.39\times 10^{-3}$ 49 3501
Table 4: [Corresponds to Figure 8] Performance metrics for the viscous Burgers equation using single-precision (32-bit) arithmetic for cases shown in Figure 8: The relative $L_{2}$ error, training time, and number of training parameters are evaluated for solving the viscous Burgers equation with various optimizers and combinations of line search algorithms. The convergence criteria are based on absolute and relative tolerances in norm of the gradient of the loss function, which is set as $[\text{ATOL},\text{RTOL}]=[10^{-7},10^{-7}]$. This experiments shows that Adam combined with BFGS trust-region provides better accuracy (hghlighted in bold fonts) .
Figure 9: Convergence of PINNs for the viscous Burgers equation under various combinations of line search and BFGS algorithms for double-precision arithmetic (64-bit floats) with GPU fully saturated with model parameters to avoid the effect of latency on compute time: The left panel illustrates a scenario where the iteration count of the first-order Adam algorithm dominates, and combined with BFGS using the trust-region approach. This pairing unlike Figure 8 does not exhibit degeneracy during BFGS iterations due to double-precision and converges to much better loss value. The right panel presents the convergence behavior of BFGS paired with trust-region and backtracking line search methods. Inline to the Figure 8. in this setup as well, Adam is applied for 11 iterations as a warmup and then switched to BFGS algorithms. The BFGS-based optimizer requires significantly more iterations than first-order optimization methods like Adam and Rectified Adam (RAdam) [79].
@@
Table 5: [Corresponds to Figure 9] Performance metrics for the viscous Burgers equation using double-precision (64-bit) arithmetic for cases shown in Figure 9: The relative $L_{2}$ error, training time, and number of training parameters are evaluated for solving the viscous Burgers equation with various optimizers and combinations of line search algorithms. The convergence criteria are based on absolute and relative tolerances in norm of the gradient of the loss function and number of iterations, and set as $[\text{ATOL},\text{RTOL}]=\{[10^{-8},10^{-8}]\leavevmode\nobreak\ \leavevmode\nobreak\ ||\leavevmode\nobreak\ \leavevmode\nobreak\ (\#\leavevmode\nobreak\ \leavevmode\nobreak\ \text{of steps}=30000$).}. This experiments shows that RAdam combined with BFGS with backtarcking linesearch provides better accuracy (hghlighted in bold fonts). The computation is performed on Nvidia-GPU Card RTX-3090 with persistence memory usage of 76 % (Total 30 GB) and compute usage of 99% i.e. 550.44 GFLOPs of total theoretical peak of 556 GFLOPs.
Figure 10: Convergence of PINNs for the viscous Burgers equation under various combinations of line search and BFGS algorithms for single-precision arithmetic (32-bit floats) with the GPU fully saturated by collocation points and data points (initial and boundary conditions) to mitigate the impact of latency on computation time: The left panel depicts a scenario where the iteration count for the first-order Adam and RAdam optimizers is fixed at 1001, followed by the second-order BFGS optimizer using the trust-region and backtracking line search algorithms. Similar to the cases shown in Figure 8, this setup also exhibits degeneracy during BFGS iterations due to limited numerical precision. The right panel shows the convergence behavior of the standalone BFGS optimizer with trust-region and backtracking line search methods, without any warmup by a first-order optimizer. In this configuration, the loss stagnates at $10^{-4}$ due to insufficient precision, particularly for the Hessian matrix inversion required for descent.
Adam [1001] + BFGS with trust-region [2142] $1.51\times 10^{-3}$ 30 1341
Adam [1001] + BFGS with backtracking [4544] $1.41\times 10^{-3}$ 43 1341
radam [1001] + BFGS with backtracking [4565] $1.43\times 10^{-3}$ 41 1341
BFGS with backtracking [2013] $1.65\times 10^{-3}$ 18 1341
BFGS with trust-region [2039] $\mathbf{1.1}\times\mathbf{10^{-3}}$ 24 1341
Table 6: [Corresponding to Figure 10] Performance metrics for the viscous Burgers equation using single-precision (32-bit) arithmetic for the cases shown in Figure 10: The relative $L_{2}$ error, training time, and number of training parameters are reported for solving the viscous Burgers equation with various optimizers and combinations of line search algorithms. The convergence criteria are defined by absolute and relative tolerances on the gradient norm of the loss function, set as $[\text{ATOL},\text{RTOL}]=[10^{-7},10^{-7}]$. The results demonstrate that BFGS with the trust-region line search achieves the highest accuracy (highlighted in bold) among all cases presented in Figure 10.
Figure 11: Convergence of PINNs for the viscous Burgers equation under various combinations of line search and BFGS algorithms for double-precision arithmetic (64-bit floats) with the GPU fully saturated by collocation points and data points (initial and boundary conditions) to minimize the impact of latency on computation time: The left panel illustrates a case where the iteration count for the first-order Adam and RAdam optimizers is fixed at 1001. Unlike Figure 10, this combination does not exhibit degeneracy during BFGS iterations due to the use of double-precision arithmetic and achieves a significantly lower loss value. The right panel presents the convergence behavior of the standalone BFGS optimizer with trust-region and backtracking line search methods, without warmup from a first-order optimizer. Notably, algorithms employing the backtracking line search method demonstrate better convergence rates and higher accuracy.
@@
Table 7: [Corresponding to Figure 11] Performance metrics for the viscous Burgers equation using single-precision (64-bit) arithmetic for the cases shown in Figure 11: The relative $L_{2}$ error, training time, and number of training parameters are reported for solving the viscous Burgers equation with various optimizers and combinations of line search algorithms. The convergence criteria here is based on union of absolute and relative tolerance of norm of gradient and considered as: $[\text{ATOL},\text{RTOL}]=\{[10^{-7},1\times 10^{-7}]\leavevmode\nobreak\ \leavevmode\nobreak\ ||\leavevmode\nobreak\ \leavevmode\nobreak\ (\#\leavevmode\nobreak\ \leavevmode\nobreak\ \text{of steps}=30000$).}. The results demonstrate that RAdam paired with BFGS with the backtracking line search achieves the highest accuracy (highlighted in bold) among all cases presented in Figure 11. The computation is performed on Nvidia-GPU Card RTX-3090 with persistence memory usage of 75 % (Total 24.56 GB) and compute usage of 99% i.e. 429 GFLOPs of total theoretical peak of 433.9 GFLOPs.
@@
To customize an instance of the BFGS optimizer, the choices for items (1)–(5) should be made based on the specific problem being addressed. For this purpose, we employed the state-based Optimistix library [84], which integrates seamlessly with the Optax library, to implement various variants of the BFGS optimizer by inheriting the AbstractBFGS class. These variants are tailored according to the selected line search algorithms, tolerance levels, convergence criteria, and types of Newton descent methods. For example, the BFGS optimizer with a trust-region line search algorithm and $L_{\infty}$-norm convergence is shown in code listing Figure 7. To evaluate the efficiency of the optimizers in terms of both accuracy and runtime on a GPU, we performed computational experiments in two distinct scenarios. In the first scenario, GPU saturation is ensured by setting the model parameters to a sufficiently large value, thus avoiding latency from impacting compute time. In the second scenario, saturation is achieved by increasing the number of collocation points.
In Scenario 1, the convergence history for single-precision (32-bit float) arithmetic is presented in Figure 8. This convergence is achieved using a neural network with 8 hidden layers, each containing 20 neurons, a $\tanh$ activation function, 200 randomly sampled data points for the initial and boundary conditions, and 10,000 collocation points to calculate the residual loss. The performance of five different optimizer combinations is summarized in Table 4. The left panel of Figure 8 shows a scenario where the iteration count of the first-order Adam algorithm dominates, combined with BFGS using the trust-region line search. The key observation from this setup is that when Adam is run for an extended period, the BFGS optimizer provides no additional advantages due to the degeneration of the loss function caused by single-precision floating-point representation [Row 1 of Table 4]. However, when Adam or RAdam is used with a small number of iterations (11 in this case) as a warm-up followed by BFGS with both line search algorithms, the error improves by an order of magnitude [Rows 2, 3, and 4 of Table 4]. In contrast, starting directly with the BFGS optimizer results in errors of a similar magnitude to those in the first scenario (left panel of 8). All runs in Table 4 are terminated once the absolute and relative tolerances of $[\text{ATOL},\text{RTOL}]=[10^{-7},10^{-7}]$ for the $L_{\infty}$ norm of the gradient of the loss function are met. Consequently, the number of iterations reported for the BFGS optimizers corresponds to the step where the specified tolerance is reached. The runtime for each case is also recorded in Table 4. For all runs, the runtime is less than one minute, except for the run where Adam is used for a larger number of iterations Figure (8).
To address the issue of loss degeneration and demonstrate that it is caused by precision limitations, all cases presented in Figure 8 and Table 4 are rerun using double-precision arithmetic (64 bits float) but with same hyperparameters used for 8. The convergence criteria are based on absolute and relative tolerances in norm of the gradient of the loss function and number of iterations, and set as $[\text{ATOL},\text{RTOL}]=\{[10^{-8},10^{-8}]\leavevmode\nobreak\ \leavevmode\nobreak\ ||\leavevmode\nobreak\ \leavevmode\nobreak\ (\#\leavevmode\nobreak\ \leavevmode\nobreak\ \text{of steps}=30000$)}. The convergence plot with Adam having 16001 iterations and then using BFGS paired with trust-region linesearch algorithm is shown in left subfigure of Figure 9. Unlike single precision, the BFGS method achieves better accuracy in this case, converging to an $L_{2}$ error of $2.8\times 10^{-5}$. Convergence results for other scenarios, similar to those in Figure 8, are displayed in the right subfigure of Figure 9. Performance metrics for these runs are provided in Table 5. Notably, the degeneration of the loss function is not observed and no longer impacts the convergence of the BFGS optimizer for any of the runs. Furthermore, relative $L_{2}$ errors for all the cases are reduced by two orders of magnitude compared to single-precision performance. For the remaining cases, the relative $L_{2}$ error is reduced to a minimum of $2\times 10^{-6}$. However, runs in Table 5 with 30,000 BFGS iterations indicate that the optimizer failed to converge, likely due to the increased sensitivity of the tolerance limit in double precision.
@@
In Scenario 2, where the GPU is fully utilized with collocation points and the PINN employs a smaller neural network, the convergence history for single and double precision arithmetic is shown in Figures 10 and 11, respectively. Convergence was achieved using a neural network with four hidden layers, each containing 20 neurons, a $\tanh$ activation function, 250 randomly sampled data points for initial and boundary conditions, and 50,000 collocation points. The performance metrics, including the relative $L_{2}$ error and training time for single and double precision, are summarized in Tables 6 and 7, respectively. The interpretation of these results is similar to that of Scenario 1, though the overall accuracy is slightly reduced. However, in double precision, the BFGS optimizer failed to converge within 30,000 iterations, likely due to the limited representational capacity of the neural network.
@@ 2.3 Allen-Cahn equation
where $\epsilon=10^{-4}$, $\kappa=5$, and the spatio-temporoal domain for computing the solution is $[x,\leavevmode\nobreak\ t]\in[-1,1]\times[0,1]$. The periodicity at the boundaries is enforced using interpolation polynomials $\{v^{(i)}(x)\}_{i=1}^{n}$ (see [85]) and defined as
@@
Table 10: Double-Precision PINNs for the Allen-Cahn equation: Case 2: A PINN with three layers, each containing 30 neurons, is trained for 1000 iterations using the Adam optimizer, followed by 20,000 iterations with BFGS, SSBroyden, and L-BFGS optimizers with Wolfe line-search. Case 3: The same network structure as Case 2 is used, but the number of iterations with BFGS and SSBroyden is reduced to 30,000. Case 4: The same network structure as in Case 2 and Case 3 is used; however, instead of starting with the Adam optimizer, the training begins directly with second-order optimizers, namely BFGS and SSBroyden.
@@
6 Adam [5000] + BFGS with Wolfe-Bspline $1.17\times 10^{-2}$ 11018.23 5,540
@@ 2.4 Kuramoto-Sivashinsky equation
Figures 16 and 17 present the results for the Kuramoto-Sivashinsky equation for Case 2. Figure 16 compares the performance of BFGS and SSBroyden in predicting the Kuramoto-Sivashinsky equation using double-precision PINNs with 20 time windows, displayed through contour plots. The absolute error for each prediction is shown, highlighting a significant difference in error magnitude between the two optimizers, particularly around $t=1.0$, where BFGS struggles to accurately predict the solution. Figure 17 illustrates the performance of BFGS and SSBroyden at three time steps: $t=0.0$, $t=0.5$, and $t=1.0$ across the spatial domain for Case 2. The figure demonstrates the complexity of the Kuramoto-Sivashinsky equation solution as it evolves over time. Although predicting the solution around $t=1.0$ is challenging, the network successfully predicts the solution, with the prediction and the true solution matching closely.
@@ 2.5 Ginzburg-Landau equation
Figure 26: Double-precision PINNs for the Stokes equation: The convergence of Physics-Informed Neural Networks (PINNs) for the lid-driven wedge flow governed by the Stokes Equation 23 is studied using the Adam optimizer and SSBFGS, combined with trust-region and backtracking line search methods. All computations are performed using 64-bit floating-point arithmetic. The convergence results indicate that combining SSBFGS with the trust-region line search method significantly outperforms the other two approaches.
#Item Optimizer[# Iters.], Line-search [#Iters.] Relative $l_{2}$ error (u, v, p) Total parameters
1 Adam (18400) + SS-BFGS with Trustregion (419965) ($3.25\times 10^{-4}$, $1.57\times 10^{-4}$, $1.1\times 10^{-1}$) 33,667
2 Adam (18400) + SS-BFGS with backtracking (178741) ($3.26\times 10^{-4}$, $1.60\times 10^{-4}$, $1.1\times 10^{-1}$) 33,667
3 Adam (500000) ($2.7\times 10^{-2}$, $2.58\times 10^{-2}$, $1.1\times 10^{-1}$) 33,667
Table 16: Double-precision PINNs for the Stokes equation: Comparison of relative $l_{2}$ errors, and total number of parameters for PINNs using SSBFGS (with trustregion and backtracking line search) and Adam optimizes. we train the model for all the three cases using same number of model parameters. The absolute and relative tolerance for convergence of SSBFGS optimizers is set $10^{-14}$ and $10^{-14}$.The computation is performed on Nvidia-GPU Card H100 with persistence memory usage of 75% (Total 80 GB) and compute usage of 99%, i.e. 25.35 TFLOPs of total theoretical peak of 25.61 TFLOPs.
@@ 2.6 Lid-driven wedge flow
[…]
@@
The boundary conditions and domain configuration are illustrated in Figure 27. This problem is primarily considered to assess the effectiveness of the proposed optimizer for solving PDEs in scenarios involving extremely small values. Moffatt [90] derived an asymptotic solution to Equation 23 using a similarity approach, demonstrating that the strength of eddies depends on the wedge angle. For the specific case depicted in Figure 27, which features a wedge angle of 25.53, the strength of each successive eddy should asymptotically be about 400 times weaker than the preceding one. Accurately resolving such weak eddies requires a highly precise and robust optimizer. This problem has not yet been addressed using the PINNs presented in previous studies, primarily due to two key challenges: the limitations of first-order optimizers and the difficulty of efficiently performing double-precision computations. Even in conventional numerical method these weaker eddies are captured with high-order numerical methods [91]. In this work, we conduct computational experiments to overcome these challenges by comparing our implementation of SSBFGS employing both backtracking and trustregion line search methods—with the commonly used Adam optimizer. Therefore, this example also answers the long standing question of PINN recovering the solution to machine precision. The results of this comparison are discussed in the following paragraphs.
@@
In Figure 27, we show the fluid domain and boundary condition on each edge of the domain with non zero transverse velocity $(u)$ at the top edge and rigid wall (no-slip) condition on the other two edges. To validate the solution obtained by PINNs we compute the reference solution by solving Equation 23 using spectral element method as proposed in [91] and use the Nektar++ codebase [92] to perform the simulation. We show the True streamlines in Figure 27b obtained from the Nektar++. To configure the PINN across all cases, we employ a neural network architecture consisting of 8 fully connected layers, each with 64 neurons, and use the $\tanh$ activation function. A total of 120,000 residual points are sampled uniformly from the domain using $\mathcal{U}[0,1)$. To enhance convergence speed and accuracy, we adopt residual-based attention as proposed in [32], along with the hard imposition of boundary conditions following [93]. Training is conducted for three distinct cases. In the first case, the Adam optimizer is used. In the second and third cases, we employ the SSBFGS optimizer in combination with trust-region and backtracking line search strategies, respectively. For the second-order optimization in these cases, the training is initially warmed up with Adam for 18,401 iterations before switching to SSBFGS. The convergence history for all three cases is presented in Figure 26. As shown in Figure 26, the combination of SSBFGS with the trust-region line search demonstrates superior performance compared to both SSBFGS with backtracking and the Adam optimizer. Specifically, SSBFGS with trust-region converges to a loss value of $10^{-12}$, while SSBFGS with backtracking and the Adam optimizer converge to $10^{-10}$ and $10^{-5}$, respectively.
@@
Training details for all three cases are summarized in Table 16. To evaluate the accuracy of each optimizer, we visualize the resulting streamlines in Figure 28. As illustrated (Figure 28a), the combination of SSBFGS with the trust-region line search successfully recovers all four eddies, including the weakest one with a velocity magnitude of $|u|=10^{-8}$. In contrast, SSBFGS with backtracking (Figure 28b) captures only the dominant eddy accurately, failing to resolve the weaker structures. The streamlines produced using the Adam optimizer (Figure 28c) do not accurately capture any eddies. It is worth noting that although both SSBFGS variants yield similar relative $L_{2}$ errors, this metric reflects the mean error and does not necessarily indicate uniform convergence. Consequently, the streamlines generated by SSBFGS with backtracking significantly deviate, as streamlines are highly sensitive to the precision of the velocity components $(u,v)$. Finally, in Figure 28d, we present the Moffatt eddy strength [90], which compares the centerline transverse velocity magnitude $(|u|)$—indicated by the dashed line in Figure 27(b)—as extracted from Figure 28a, Figure 28b, and Figure 28c. It is evident that the SSBFGS optimizer with trust-region line search most accurately captures the weakest eddies, outperforming both the backtracking variant and the Adam optimizer.
@@
(a) SS-BFGS Trustregion
(b) SS-BFGS backtracking
@@
Figure 28: Solutions of Equation 23 obtained using PINN optimized with (a) SSBFGS optimizer with trustregian based linesearch and subfigure (b) represents the streamlines obtained by pairing the SSBFGS with backtracking linesearch. (c) represents the streamlines obtained while using Adm optimizer. Subfigure (d) shows comaprison of centre-line (shown with dashed line in Figure 27(b) transverse velocity $(|u|)$ in (a), (b) and (c) also known as Moffatt eddy strength.

### v5

@@ 2.1.2 Performance of optimzers based on composition operator in the JAX ecosystem
Figure 5: Scenario 1: Double-Precision PINNs for the Burgers Equation. Results are shown with backtracking and trust-region strategies. The GPU is fully saturated with model parameters to avoid the effect of latency on compute time. The left panel illustrates a scenario where the iteration count of the first-order Adam algorithm dominates, and combined with BFGS using the trust-region approach. This pairing unlike Figure 29 does not exhibit degeneracy during BFGS iterations due to double-precision and converges to much better loss value. The right panel presents the convergence behavior of BFGS paired with trust-region and backtracking line search methods. Inline to the Figure 29. in this setup as well, Adam is applied for 11 iterations as a warmup and then switched to BFGS algorithms. The BFGS-based optimizer requires significantly more iterations than first-order optimization methods like Adam and Rectified Adam (RAdam) [79].
@@
[…]
Figure 6: Scenario 2: Double-Precision PINNs for the Burgers Equation. Results are shown with backtracking and trust-region strategies. The GPU is fully saturated by collocation points and data points (initial and boundary conditions) to minimize the impact of latency on computation time. The left panel illustrates a case where the iteration count for the first-order Adam and RAdam optimizers is fixed at 1001. Unlike Figure 30, this combination does not exhibit degeneracy during BFGS iterations due to the use of double-precision arithmetic and achieves a significantly lower loss value. The right panel presents the convergence behavior of the standalone BFGS optimizer with trust-region and backtracking line-search methods, without warmup from a first-order optimizer. Notably, algorithms employing the backtracking line-search method demonstrate better convergence rates and higher accuracy.
@@
[…]
@@
To customize an instance of the BFGS optimizer, the choices for items (1)–(5) should be made based on the specific problem being addressed. For this purpose, we employed the state-based Optimistix library [84], which integrates seamlessly with the Optax library, to implement various variants of the BFGS optimizer by inheriting the AbstractBFGS class. These variants are tailored according to the selected line search algorithms, tolerance levels, convergence criteria, and types of Newton descent methods. For example, the BFGS optimizer with a trust-region line search algorithm and $L_{\infty}$-norm convergence is shown in code listing Figure 4. To evaluate the efficiency of the optimizers in terms of both accuracy and runtime on a GPU, we conducted computational experiments under two distinct settings. In Scenario 1, GPU saturation is achieved by increasing the number of model parameters to a sufficiently large value, thereby minimizing the impact of latency on compute time. In Scenario 2, saturation is obtained by increasing the number of collocation points while keeping the model size fixed.
[…]
@@
In Scenario 2, where the GPU is fully utilized with collocation points and the PINN employs a smaller neural network, the convergence history for double precision arithmetic is shown in Figure 6. Convergence was achieved using a neural network with four hidden layers, each containing 20 neurons, a $\tanh$ activation function, 250 randomly sampled data points for initial and boundary conditions, and 50,000 collocation points. The performance metrics, including the relative $L_{2}$ error and training time are summarized in Table 5. The interpretation of these results is similar to that of Scenario 1, though the overall accuracy is slightly reduced. However, in double precision, the BFGS optimizer failed to converge within 30,000 iterations, likely due to the limited representational capacity of the neural network.
@@ 2.3 Allen-Cahn equation
where $\epsilon=10^{-4}$, $\kappa=5$, and the spatio-temporoal domain for computing the solution is $x,\leavevmode\nobreak\ t]\in[-1,1]\times[0,1]$. The periodicity at the boundaries is enforced using interpolation polynomials $\{v^{(i)}(x)\}_{i=1}^{n}$ (see [85]) and defined as
@@
Table 8: Double-Precision PINNs for the Allen-Cahn equation: Case 2: A PINN with three layers, each containing 30 neurons, is trained for 1000 iterations using the Adam optimizer, followed by 20,000 iterations with BFGS, SSBroyden, and L-BFGS optimizers with Wolfe line-search. Case 3: The same network structure as Case 2 is used, but the number of iterations with BFGS and SSBroyden is reduced to 30,000. Case 4: The same network structure as in Case 2 and Case 3 is used; however, instead of starting with the Adam optimizer, the training begins directly with quasi-Newton optimizers, namely BFGS and SSBroyden.
@@
6 Adam [5000] + BFGS with Wolfe-Bspline $17\times 10^{-2} 11$018.23 5,540
@@ 2.4 Kuramoto-Sivashinsky equation
Figures 11 and 12 present the results for the Kuramoto-Sivashinsky equation for Case 2. Figure 11 compares the performance of BFGS and SSBroyden in predicting the Kuramoto-Sivashinsky equation using double-precision PINNs with 20 time windows, displayed through contour plots. The absolute error for each prediction is shown, highlighting a significant difference in error magnitude between the two optimizers, particularly around $t=1.0$, where BFGS struggles to accurately predict the solution. Figure 12 illustrates the performance of BFGS and SSBroyden at three time steps: $t=0.0$, $t=0.5$, and $t=1.0$ across the spatial domain for Case 2. The figure demonstrates the complexity of the Kuramoto-Sivashinsky equation solution as it evolves over time. Although predicting the solution around $t=1.0$ is challenging, the network successfully predicts the solution, with the prediction and the true solution matching closely.
@@ 2.5 Ginzburg-Landau equation
Figure 21: Double-precision PINNs for the Stokes equation ( Equation 16). The model is trained using the Adam optimizer followed by SSBFGS with trust-region and backtracking line search strategies. All computations are performed using 64-bit floating-point arithmetic. The results show that SSBFGS combined with the trust-region method significantly outperforms other approaches in terms of convergence.
Optimizer [# Iters.], Line Search [# Iters.] Relative $l_{2}$ Error (u, v, p) Total Parameters
Adam (18400) + SSBFGS with Trust-Region (419,965) ($3.25\times 10^{-4}$, $57\times 10^{-4}, 1.1$$10\times 10^{-1}$) 33,667
Adam (18400) + SSBFGS with Backtracking (178,741) ($3.26\times 10^{-4}$, $60\times 10^{-4}, 1.1$$10\times 10^{-1}$) 33,667
Adam (500,000) ($2.70\times 10^{-2}$, $2.58\times 10^{-2}$, $1.10\times 10^{-1}$) 33,667
Lion (500,001) ($1.63$, $1.01$, $9.98$) 33,667
AdaBelief (500,001) ($3.11\times 10^{-2}$, $3.53\times 10^{-2}$, $1.50\times 10^{-1}$) 33,667
Adam (18,401) + Nonlinear CG (Polak–Ribiere) (500,001) [90] ($4.13\times 10^{-2}$, $3.96\times 10^{-2}$, $9.54\times 10^{-1}$) 33,667
Table 14: Double-precision PINNs for the Stokes equation. Comparison of relative $l_{2}$ errors and total number of model parameters for different optimizers. All models use the same network architecture with 33,667 parameters. The absolute and relative tolerance for SSBFGS convergence is set to $10^{-14}$. SSBFGS with trust-region line search achieves the best accuracy. For comparison, we also report the performance of Lion, AdaBelief, and Hessian-free Nonlinear Conjugate Gradient (CG) using the Polak–Ribiere variant [90], all of which perform significantly worse. All computations are performed on an NVIDIA H100 GPU, with persistent memory usage of 75% (total 80 GB) and 99% compute utilization, achieving 25.35 TFLOPs of the theoretical peak of 25.61 TFLOPs.
@@ 2.6 Lid-driven wedge flow
$\begin{aligned} \frac{\partial p}{\partial x}-\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right)=0,\\ \frac{\partial p}{\partial y}-\left(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\right)=0,\\ \frac{\partial u}{\partial x}+\frac{\partial v}{\partial y}=0,\end{aligned}$ (16)
@@
The boundary conditions and domain configuration are illustrated in Figure 22. This problem is primarily considered to assess the effectiveness of the proposed optimizer for solving PDEs in scenarios involving extremely small values. Moffatt [91] derived an asymptotic solution to Equation (16) using a similarity approach, demonstrating that the strength of eddies depends on the wedge angle. For the specific case depicted in Figure 22, which features a wedge angle of 25.53, the strength of each successive eddy should asymptotically be about 400 times weaker than the preceding one. Accurately resolving such weak eddies requires a highly precise and robust optimizer. This problem has not yet been addressed using the PINNs presented in previous studies, primarily due to two key challenges: the limitations of first-order optimizers and the difficulty of efficiently performing double-precision computations. Even in conventional numerical method these weaker eddies are captured with high-order numerical methods [92]. In this work, we conduct computational experiments to overcome these challenges by comparing our implementation of SSBFGS in JAX—employing both backtracking and trustregion line search methods—with the commonly used Adam optimizer. Therefore, this example also answers the long standing question of PINN recovering the solution to machine precision. The results of this comparison are discussed in the following paragraphs.
@@
In Figure 22, we show the fluid domain and boundary condition on each edge of the domain with non zero transverse velocity $(u)$ at the top edge and rigid wall (no-slip) condition on the other two edges. To validate the solution obtained by PINNs we compute the reference solution by solving Equation (16) using spectral element method as proposed in [92] and use the Nektar++ codebase [93] to perform the simulation. We show the True streamlines in Figure 22b obtained from the Nektar++. To configure the PINN across all cases, we employ a neural network architecture consisting of 8 fully connected layers, each with 64 neurons, and use the $\tanh$ activation function. A total of 120,000 residual points are sampled uniformly from the domain using $\mathcal{U}[0,1)$. To enhance convergence speed and accuracy, we adopt residual-based attention as proposed in [32], along with the hard imposition of boundary conditions following [94]. Training is conducted for three distinct cases. In the first case, the Adam optimizer is used. In the second and third cases, we employ the SSBFGS optimizer in combination with trust-region and backtracking line search strategies, respectively. For the second-order optimization in these cases, the training is initially warmed up with Adam for 18,401 iterations before switching to SSBFGS. The convergence history for all three cases is presented in Figure 21. As shown in Figure 21, the combination of SSBFGS with the trust-region line search demonstrates superior performance compared to both SSBFGS with backtracking and the Adam optimizer. Specifically, SSBFGS with trust-region converges to a loss value of $10^{-12}$, while SSBFGS with backtracking and the Adam optimizer converge to $10^{-10}$ and $10^{-5}$, respectively. Training details for all three cases are summarized in Table 14. To further evaluate the optimizer’s performance, we also include comparison metrics for several niche first-order optimizers—Lion [95], AdaBelief [96], and the non-linear conjugate gradient method with the Polak–Ribiere update rule [90]—using the same time budget, also presented in Table 14. Notably, the SSBFGS optimizer with a trust-region approach outperforms all other methods.
@@
To evaluate the accuracy of each optimizer, we visualize the resulting streamlines in Figure 23. As illustrated (Figure 23a), the combination of SSBFGS with the trust-region line search successfully recovers all four eddies, including the weakest one with a velocity magnitude of $|u|=10^{-8}$. In contrast, SSBFGS with backtracking (Figure 23b) captures only the dominant eddy accurately, failing to resolve the weaker structures. The streamlines produced using the Adam optimizer (Figure 23c) do not accurately capture any eddies. It is worth noting that although both SSBFGS variants yield similar relative $L_{2}$ errors, this metric reflects the mean error and does not necessarily indicate uniform convergence. Consequently, the streamlines generated by SSBFGS with backtracking significantly deviate, as streamlines are highly sensitive to the precision of the velocity components $(u,v)$. Finally, in Figure 23d, we present the Moffatt eddy strength [91], which compares the centerline transverse velocity magnitude (|u|)(|u|)—indicated by the dashed line in Figure 22(b)—as extracted from Figure 23a, Figure 23b, and Figure 23c. It is evident that the SSBFGS optimizer with trust-region line search most accurately captures the weakest eddies, outperforming both the backtracking variant and the Adam optimizer.
@@
(a) $S$-BFGS Trustregion
(b) $S$-BFGS backtracking
@@
Figure 23: Solutions of Equation 16 obtained using PINN optimized with (a) SSBFGS optimizer with trustregian based linesearch and subfigure (b) represents the streamlines obtained by pairing the SSBFGS with backtracking linesearch. (c) represents the streamlines obtained while using Adm optimizer. Subfigure (d) shows comaprison of centre-line (shown with dashed line in 22(b) transverse velocity $(|u|)$ in (a), (b) and (c) also known as Moffatt eddy strength.